Find The slope of the lines that passes through the pair of points (1,7),(10,1)
Answer by mananth(16949)

x1 y1 x2 y2
1 7 10 1

slope m = (y2-y1)/(x2-x1)
( 1 - 7 )/( 10 - 1 )
( -6 / 9 )
m= - 2/3 - 2/ 3

Plug value of the slope and point ( -1 , 8 ) in
Y = m x + b
8.00 = 2/3 + b
b= 8 - 2/3
b= 7 1/ 3
So the equation will be
Y = - 2/3 x + 22/3
m.ananth@hotmail.ca